OpenStudy (animelover91804):

Use the distributive property to solve for the unknown value: 3(a+ 7) = 3* (-1) + 3 * 7

OpenStudy (anonymous):

3a+21= -3+3x7 3a+21= -3+21 3a+21= 18 -21 -21 3a=-3 /3 /3 a= -1 I think this is right......

OpenStudy (animelover91804):

♥ Thanks chu!!!!!!!!! ♥